The marginal cost of a trader has been found to be MC = 3Q2 + 8Q + 400 . Determine the total variable cost of producing 100 units of the trade’s product. 


1
Expert's answer
2021-12-16T11:13:04-0500

Integrating the function:

"Q^3+4Q^2+400Q=100^3+4\times100^2+400\times100=1 080 0000"